# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v1

@@ 2.2 Measure-theoretic $\epsilon$-entropy
where $\{\phi_{k}\}$ is a countable dense subset of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in sense of metric $D$ is always less than $1$, and for fixed $\mu$ and $\nu$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. It follows from the weak-topology on $M(X)$ that the natural projection from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in sense of weak-topology.
Given an invariant measure $\mu\in M_{f}(X)$, we say a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\bigcap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\bigcup_{i=1}^{k}C_{i})=0$. The partition entropy of $\xi$ w.r.t. $\mu$ is defined by
@@
(cf.[GS21, Theorem 3.1, Remark 3.6], [S22, Proposition 7.3]) Let $(X,f)$ be a t.d.s. with metric $d$. Then for any $\delta\in(0,1)$,
@@ 2.3 Specification property
One says that $f$ has specification property if for any $\epsilon>0$, there is an integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists $x\in X$ such that
@@ 3 Proofs of main results
This section aims to prove the main results of this paper. Essentially, the difficult part is the estimation of lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$. We shall construct a Moran fractal set containing in $G_{K}^{\omega,U}$, which is a limit set with geometric construction. Then such set is suitable for the purpose of the computation of metric mean dimensions of saturated sets.
@@ 3.1 Packing metric mean dimension of saturated sets
We prove Theorem 1.1. To arrive this goal, we divide the proof ot Theorem 1.1 into two parts.
3.1.1 Upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$
We adapt a general method presented in [PS07, Theorem 4.1] [ZCC12, Theorem 1.1 $(2)$] and [JCZ22, Proposition 4] to obtain the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$.
@@ 3.1.1 Upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$
[Y23, Lemma 3.1] Let $(X,f)$ be a dynamical system and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$ and $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
@@
$(1)$ Let $K\subset M_{f}(X)$ be a nonempty closed convex subset. Then for any $\epsilon>0$, we have
$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$
@@
$\mathcal{R}(K,\delta,n):=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
and $\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(K,\delta,n,\epsilon),$ where $S(K,\delta,n,\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{R}(K,\delta,n)$. Then by the relation $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}(K,\delta,n)$ holds for every $\delta$, it follows from Proposition 2.1 that
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}(K,\delta,n),\epsilon)\leq\sup_{k}h_{top}(f,\mathcal{R}(K,\delta,n),\epsilon).$
@@
Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}(K,\delta_{k},n_{k})$, and set
$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x).$
@@
$\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\geq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log S(K,\delta_{k_{j}},n_{k_{j}},\epsilon)\geq\Theta(K,\epsilon)-\gamma.$
@@
Thus, by the Lemma 3.2 we obtain an upper bound of the packing metric mean dimension of saturated set:
@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
Now set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with lexicographic order. Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$. For $(k,j)\in\mathcal{A}$, we shall use $(k,j)^{\ast}$ and $(k,j)_{\ast}$ to represent the one before and the one after $(k,j)$, respectively. Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall write $\delta_{k}:=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is a positive integer given in the definition of specification property.
For any non-empty open set $U$, there exist $\tilde{\epsilon}>0$, a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$. One may assume that $\epsilon^{\ast}<\tilde{\epsilon}$. Set
@@
Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. The following measure approximation derived from a slight modification of [LL24, Lemma 2.1], so we omit the proof.
@@
#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right].$
@@
For $n\in\mathbb{N}$, $\varphi\in C(X)$, write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X,[0,1])$,
@@
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Now suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how construct $\mathcal{I}_{(k,j)}$. Let
@@
$\mathbb{F}\subset G_{K}^{\omega,U}=G_{K}\cap U\cap\{x\in X:X\subset\omega_{f}(x)\}.$
@@
Firstly, we know $\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset\overline{B(a_{0},\tilde{\epsilon})}\subset U$. For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$. Hence for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there are some $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$, which implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have $\mathbb{F}\subset\{x\in X:X\subset\omega_{f}(x)\}$.
@@
Case $(1)$. If there exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
@@
Case $(2)$. If there exists $(k,j)\in\mathcal{A}$ with $j\neq 0$, such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
@@
Case $(3)$. If there exists $(k,1)\in\mathcal{A}$ and $i>1$, such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
@@
Case $(4)$. If there exists $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$, such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
@@
For simplicity, we write $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$ instead of $\mathcal{I}_{k}$, $\mathbb{F}_{k}$, respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$, which implies that $\rho(\mathbb{F})=1$. By (3.9), we have

### v2

@@ 2.2 Measure-theoretic $\epsilon$-entropy
where $\{\phi_{k}\}$ is a family of countable dense subset of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in sense of metric $D$ is always less than $1$, and for fixed $\mu$ and $\nu$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. It follows from the weak-topology on $M(X)$ that the natural projection from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in sense of weak-topology.
Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\bigcap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\bigcup_{i=1}^{k}C_{i})=0$. The partition entropy of $\xi$ w.r.t. $\mu$ is defined by
@@
(cf.[GS21, Theorem 3.1, Remark 3.6], [S22, Proposition 7.3]) Let $(X,f)$ be a t.d.s. with a metric $d$. Then for any $\delta\in(0,1)$,
@@ 2.3 Specification property
One says that a t.d.s. $(X,f)$ has specification property if for any $\epsilon>0$, there is an integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists $x\in X$ such that
@@ 3 Proofs of main results
This section aims to prove the main results of this paper. Essentially, the difficult part is the estimation of lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$. We shall construct a Moran fractal set containing in $G_{K}^{\omega,U}$, which is a limit set with geometric construction. Then such set is suitable for the purpose of the computation of the lower bound of metric mean dimensions of saturated sets.
@@ 3.1 Packing metric mean dimension of saturated sets
We first prove Theorem 1.1. To arrive this goal, we divide the proof of the Theorem 1.1 into two parts.
3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
We adapt a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to obtain the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$.
@@ 3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
[Y23, Lemma 3.1] Let $(X,f)$ be a dynamical system and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
@@
$(1)$ Let $K\subset M_{f}(X)$ be a nonempty closed convex subset. Then for any $\epsilon>0$,
$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)~{}\text{for infinitely many }n>k\right\},$
$\mathcal{Q}_{K}(\delta,n):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
and $\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(K,\delta,n,\epsilon),$ where $S(K,\delta,n,\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$. Fix $x\in G^{K}$. Then, we can choose a subsequence $\{n_{i}\}$ such that $\mathcal{E}_{n_{i}}(x)\to\mu$ for some $\mu\in K$. This implies for every $\delta>0$, there exists a positive integer $k$ such that $\mathcal{E}_{n_{i}}(x)\in B(K,\delta)$ for sufficiently large $n_{i}$ greater than $k$. Then by the relation $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ holds for every $\delta$, it follows from Proposition 2.1 that
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon),$
which implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon).$
@@
Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$, and set
$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x)\in B(K,\delta_{k}).$
@@
$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log S(K,\delta_{k_{j}},n_{k_{j}},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
Thus, by the Lemma 3.2 we obtain an upper bound of the packing metric mean dimension of $G_{K}$:
@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
Now the set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with lexicographic order. Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$. For $(k,j)\in\mathcal{A}$, we shall use $(k,j)^{\ast}$ and $(k,j)_{\ast}$ to represent the one before and the one after $(k,j)$, respectively. Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall write $\delta_{k}:=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is a positive integer given in the definition of specification property.
For any nonempty open set $U$, there exist $\tilde{\epsilon}>0$ and a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$. One may assume that $\epsilon^{\ast}<\tilde{\epsilon}$. Set
@@
Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. The following measure approximation of $\mu^{*}$ derived from a slight modification of [LL24, Lemma 2.1], so we omit the proof.
@@
#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right].$
@@
For $n\in\mathbb{N}$, $\varphi\in C(X)$, write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X,[0,1])$ taking values in $[0,1]$,
@@
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Now suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how to construct $\mathcal{I}_{(k,j)}$. Let
@@
$\mathbb{F}\subset G_{K}^{\omega,U}=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\}.$
@@
Firstly, we know $\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset\overline{B(a_{0},\tilde{\epsilon})}\subset U$. For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$. Hence for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there is a $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$, which implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have $\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}$.
@@
Case $(1)$. There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
@@
Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$, such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
@@
Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$, such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
@@
Case $(4)$. There exists $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$, such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
@@
For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$, respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$, which implies that $\rho(\mathbb{F})=1$. By (3.9), we have